Synchromania

Rick van Rein

April 24, 2000

“C ome on, father, we have to keep up with developments in the new world!” Mekwa-Nuh sighed. He wished he had never sent his son to North America to obtain a college degree. He had thought it would be best for their South African tribe. He could feel his life was not going to last much longer, and then his son would follow him up in leading the tribe. Six years ago, he had foreseen that his son needed a thorough education, so that the decisions he would have to make as a leader would be founded in reason.

What he hadn’t foreseen though, was that the western culture of America would have such an impact on his son. “Father, you didn’t send me out to study aesthetics and do nothing with it, right?” “Right, son, but...” “Well father... dear, old, wise Mekwa-Nuh, this is a place for improvement. If we can get the drums of all South African tribes to play synchronously, then we would not be displeased by the drums of neighbouring tribes interfering with our music anymore. It would all be so much more pleasing, so natural...” And his father sighed deeply, realizing that his son would make this plan come true, either now with his blessings or after he had given the tribe leadership to his son. “Son, if you think this is the way to go, then so it shall be.”

* * * * *

T he girl with the long, blond hair looked up. Jay had never noticed how much those deep blue eyes worked on him. She stood up and walked towards him, smiling pleasantly. He almost forgot the sound of the seagulls and the feeling of the tide, now rushing in, pounding against his ankles. As she embraced him, Jay wondered if her lips would taste as salty as the water of the sea, but he never got a chance to find out. At the instant she touched him, he heard drums playing in the woods. A steady, boring bash on some low-pitched drum, played without too much fantasy. Despair showed up in her heavenly blue eyes, and she whispered to him “Can you save me?” Jay did not know how. He felt the wind blowing hard and shaking him from left to right. Now the girl asked him “What are you thinking of? Wake up...”

“Jay, wake up, it’s 6:45, time to get up!” Jay opened his eyes, and stared into two friendly eyes, brown but real this time. He realized with dismay that the girl in his dreams was gone, but he was waken up by Mary, his beloved girl friend, which was at least as pleasant. “Good morning, Jay, what were you dreaming of?” “Oh (yawn), nothing special. I don’t know anymore. Oh... dear... do we have to listen to that noise?” Only now Jay realized that the alarm clock was playing one of those horrible Woop-Aloop tunes. “Well, Jay, I happen to like it. And it’s where times are going, with or without you.” “Well, since I don’t seem to have a choice, make it happen without me, okay?” While Mary started making the bed, he dragged himself to the shower. The hot water finally managed to wake him up fully.

As always when they drove to work together, they were discussing something. Today, the topic was Woop-Aloop. “You know, Jay, this must be the greatest thing that ever happened in music!” “Yeah, right, that’s what they said about Elvis too, but he died early, remember?” “No, really, Woop-Aloop is so harmonious, no more fights with neighbours over music.” She had a point there, Jay knew. Woop-Aloop was some weird South-African concept of synchronized rhythm, which was exploited by radio stations in America, too. All music that was in Woop-Aloop style, and it sometimes seemed to Jay there was nothing else left, used the same beat so that whatever anyone else played, it always was in sync with your own music, so that it mixed nicely. As if that point had to be pushed even further, a convertible pulled up next to theirs, playing music loudly. Indeed, it mixed nicely with their music, there were no irritating beats from the other car interfering with their rhythms. But Jay considered it ‘one of those modern things.’

Mary had already changed the subject. “You do remember that we are going to look after motorcycles during lunch, don’t you?” Jay nodded, anticipating how frantic she would get when seeing the newest racers. These things made their difference in age of almost seven years visible. Then again, Jay had also set his mind on a motorcycle, it was a boy’s dream come true for him.

Jay dropped off Mary at her office and quickly changed the radio to a classical station while driving on to his own job. Sometimes he thought he was getting too old for this world. Strange, actually, since in his engineering job he was supposed to invent futuristic devices. “No,” he thought, “I am not getting old at all. I’m just a little more settled down than many others.”

At lunch time, Jay and Mary hurried to the motorcycle dealership. Mary, as anticipated, immediately started drooling at the sight of the new Kiwisika Hirikara, and indeed, it was quite a compelling vehicle to Jay too. “Ah, madam, sir, excellent choice if I may say so!” Obviously, the opening sentence of a sales representative who receives a percentage of his sales as a bonus. “May I point out some of the features?” As Jay had expected, the bike’s features were all advantages. Mary knew all the technical blah-blah well by now, after all the brochures she’d been reading. Most of the terms were familiar to Jay too, except for one: synchronous blinking. “What is that exactly?” he demanded. The salesman smiled his best toothpaste-commercial-smile. “Synchronous blinking is a new Japanese development. You know how they work: they observe us, Americans, and copy creatively. The synchronous blinking feature is inspired on our Woop-Aloop style of music. It actually tunes in to the radio and picks up the Woop-Aloop signals and synchronizes on it. Have you ever seen cars and bikes waiting for a light, having their blinkers on? For a while, two blinking lights blink synchronously, then they divert from each other, then they are each others opposites... enough to drive a law-abiding citizen crazy. With this new feature, however, this is all in the past: all participating vehicles will blink in a synchronized fashion, and therefore the streets will look more peaceful.” Jay was not sure whether to smile in pity or to burst in laughter. So he just nodded.

But when they were back in the car, on the way to their respective jobs, Jay did notice that several cars already blinked nicely in rhythm. And then he noticed how ugly the few non-rhythmical blinkers looked. Indeed, this did seem like a pleasant feature to have!

Mary was excited about the motorcycle, Jay could tell. She produced such a waterfall of words, that she didn’t even notice that Jay brought her somewhere else than her office. Jay went straight to the Zilker Park. Only when Jay got out, politely opened her door for her and helped her out, she started getting curious. Mary inquired what he was doing, but he just smiled serenely, knowing that her curiosity would only be strengthened that way.

Jay took her to the side of the lake, and asked her to sit down on the grass. He knelt in front of her, while fetching a small piece of paper. “Er... Mary, I am not much of a poet, but this is for you.” And he started, a little reddish in the beginning but with more power as he progressed:

Words cannot express what I feel Now that I sit in front of you and kneel You make me feel like romance More than any poem, song or dance You bring me up when I am down You cause laughter when I frown You brought happiness in my life Mary, will you be my wife?

“Oh, Jay, how romantic. And I think you are a great poet! Yes! Yes! Yes! I’d love to be your wife!” She swung herself around his neck, making him fall over. They carried on for a while, giving the swans in the lake an interesting view.

After Jay had recollected his glasses, they went to the car, but not before deciding to have a romantic dinner that night.

T he table was set for two, candle light twinkled in Mary’s eyes and the smell of a delicious lamb stew was hanging around them. Mozart played softly in the background. “Dear,” Jay said, “I am looking forward to marrying you. And, maybe, we should finally go and look for that new house we’ve been dreaming about for some while now.” Mary nodded, realizing that a motorcycle would not be an issue anymore if they decided to take that step. But, actually, she was wanting to go live somewhere outside of Austin, in a friendly little town. And she realized this was what they both really wanted, even more than a motorcycle.

* * * * *

T hey were already married for a couple of months when Jay thought back to their visit to the motorcycle shop. “Can you imagine,” he asked Mary, “that we wanted to buy a motorcycle without synchronous blinking?” Mary grinned. “Yes, that would have been a really stupid decision. It is so convenient to have everything in sync.” They were moving into their new house, and unpacking their belongings brought back memories. They had bought quite some new appliances that were all synchronized, and it was so convenient and normal to them now: the phone would ring in sync with the beats of music on the radio, the washer/dryer would never switch on or off at an unexpected moment but always when, say, an explosion occurred on the television, so that it would not cause unexpected disruptive noises throughout the house like in the old days before synchronization. Indeed, life was great with synchronized appliances.

Jay went outside, to look at the garden. The smell of new born hay touched his nose. Wonderful, to live in the country. He called his wife to come out of the house. “Mary, dear, look here. Do you think we can make a small pergola here, over the terrace? It would be so nice in spring time.” “Dear, you’re the gardener in this house, if you think you can make that, go right ahead!” She walked up to him and took his hand in hers. This marriage was the best thing she’d ever done in her whole life, she was sure. And the house made it all look even better. Jay would have the garden to spend his time on, she herself had a magnificent room for her hobbies... maybe she would even pick up painting again. Her fingers started itching at the mere thought!

“Howdy, neighbours, how are ya’?” A voice over the fence. A friendly looking, slightly fattened men’s face looked over the edge. “Oh, hello, sir, nice to meet you. I am Jay, and this is Mary, my wife.” “Howdy. Nice to meet y’all. My name is Simon, and my wife’s is Pat... PAT! Will you come over here for a sec?” Pat came, a cuddly little wife, the kind that knits for her grandchildren and probably bakes great apple pies. “Hello, dears, nice to meet you. I am Pat” the prototype grandmother announced cheerfully. Jay and Mary introduced themselves to the couple. Pat seemed to like to meet new people. “Say, we’re going to have a barbecue with some other folks of around here tonight. You must join in, really.” Jay and Mary gladly accepted. It was a nice opportunity to meet some of their neighbours. “Well, see ya’ then! Nice talking to y’all!” “Yes, nice talking to you too.” And they each went their own way.

T hat evening, the smell of hardly burnt meat touched Mary’s nose, and reminded her of the appointment they had. She reminded Jay that they really had to go now. Jay was not really in the mood to go out tonight, but he realized that he should, if he wanted to get acquainted with his neighbours. So they went. They were a little late, almost everyone had already arrived, eager as they were to meet the new kids in town. The resulting open and friendly atmosphere made Jay relax quickly. The wine did the rest, and it did not take long until Jay and Mary were in pleasant conversation with the others attending the barbecue.

Their neighbours turned out to be very kind people, just like they had seemed earlier that day. Pat had been a housewife all her life long, and indeed she promised to bake an apple pie for the newcomers. Simon was a retired colonel from the US army. Jay and he could get along well, discussing the miraculous new developments that technology had brought them, while Pat gave Mary an introduction to the gossip on everyone in the neighbourhood. While talking, Mary got to meet most of the neighbours. Jay hardly got a chance to meet anyone because he was so occupied in his discussion with the colonel.

It was already late when Jay and Mary decided to leave the party and go home. They walked over to their house hand in hand, quietly enjoying the pleasantness of the new neighbourhood they had found. They already felt at home!

“J ay, will you please hook up the washer and dryer, so I can do some laundry this afternoon? When unpacking, I found back an old dress of mine and I’d like to wear it tomorrow.” Jay looked up from the television which he was trying to get connected to the cable. Well, the TV could wait, he decided. First he would work on the devices that Mary needed.

While enjoying an aria from Carmen that he whistled along with, he connected the washer and dryer for Mary. It was a small job that he had done often before, so he was done in an hour. Mary gave him a swift kiss and started selecting laundry. Jay finished hooking up the television and left the house, to weed and plant seeds in their garden. Ah, was that the voice of colonel Simon he heard over the fence? He called, and the reddish friendly face of Simon appeared, this time covered by a cowboy hat that appeared to be the one that was used to define the gallon. “Howdy, neighbour, how are ya’?” the friendly Texan inquired. Jay just wanted to explain that the headache from the party last night had gone by now, when Mary ran out of the house, almost in panic.

“Jay, have you not connected the synchronization to the washer and the television yet? It is driving me crazy how the washer keeps switching on and off its pump at the most unexpected moments.” Jay frowned. He had not noticed the problem, but then again, he had been listening to opera. Mary had probably switched the radio to some Woop-Aloop station as soon as he had left the house, and had gone frantic by asynchronous behaviour of their devices. But Jay could not imagine how the systems would possible ignore the synchronization; they automatically filtered their synchronizing signals out of the ether!

Simon had the solution. “Didn’t y’all know? We’re on the border here, between Dallas and Austin, and they each send out different signals. Some devices pick out the Dallas signals, others prefer the Austin ones, and yet others like to tune in with local transmissions.” Jay, being a technician, knew what Simon meant, but Mary needed a little more explanation. “It’s simple, dear. Radio signals take some time to travel to the receivers. Even if Dallas and Austin would send out their signals at the same moment, they would not arrive at the same time everywhere. If you live in Austin, there is no problem, the Austin beacons are so much stronger that they will dominate anything else. Here, however, we’re in the middle between major cities, and our devices get confused. It appears we have no synchronization around here, dear. Unless either Austin or Dallas decides to increase the power of their broadcasts.” “That’ll be the day,” Simon added, “Don’t y’all know what they call us out here? We’re supposed to be the ‘Jazzers’, the outcast of the modern synchronized style of living. We do not work in any sort of beat, and them folks in them large cities can’t seem to bear that. It is what them people sees as our rhythmical improvisation that gave us our nickname of Jazzers. And believe me, they look down at us for not being synchronized!”

T hat evening, Mary was watching the Dallas news bulletin while Jay was trying to solve a crossword riddle. “Honey, what’s ‘great’ with six letters?” Jay asked. He immediately got a response: “Superb!” Jay filled it in but looked up when he realized that this was not the enthousiasm Mary normally felt for crossword puzzles. And indeed, she was excited about something on TV. “What’s so pleasant, dear?” he inquired. “Oh, Jay, remember what you said this afternoon? That the only way to get our devices to synchronize again would be if either Austin or Dallas would increase the power of their broadcasts? Well, they just announced that Dallas is going to increase theirs!” Jay was puzzled, and not by his riddle this time. “Why would they ever want to do that? We can see all TV channels we wish, can’t we?” “Oh, simple! People in the suburbs around Dallas, like us, do not like to travel to Dallas if their cars and other devices have to go from one synchronization area to another. For instance, people do not like to go from Austin to Dallas, simply because they would have to change the synchronization beats that they are so accustomed to.” Mary was obviously happy, but Jay still couldn’t understand why anyone would pay extra money for such a thing. Until Mary explained to him that the entire reason was that malls around Dallas were trying to get more people interested in visiting their stores instead of Austin’s. By increasing the zones in which people were willing to travel, they would increase their number of potential customers. Mary was excited, Jay frowned.

As the news bulletin had promised, the synchronization in their little town got stabilized to the signals of the Dallas beacon. All devices recognized the stronger beacon and tuned in to it. “Oh, Jay, don’t you just love it? I’ll never hear the washer interrupt my favourite show anymore! It makes me so happy — I feel like celebrating.” “Celebrating? Like how??” “Well... how about going out tonight, visit Dallas? They’ve got a pretty lively area there, close to I-75. Shall we go there?”

Jay still wondered how he had ever come to say yes. A discotheque! As if they were teenagers! All that was played nowadays was this Woop-Aloop rubbish. And Mary, of all people, had to enjoy it. When they got out of the car they went into a nice-looking nightclub. Jay wondered if there were mechanics at work with power drills, or if it was truly music that was being played. “Come on Jay, let’s dance.” Jay followed her to the dance floor, but there were things he had done with more pleasure...

They danced all night. Mary was over-excited. “Imagine Jay, no headaches from changes in rhythm anymore if we go home in a while. Isn’t it wonderful?” “Hmm... the change won’t cause a headache for sure. But the beers might... Anyway, it was kind of nice to do for a change,” even Jay had to admit. After his first few beers, the rhythm had gotten a hold of him too.

When they went to their car, they saw other people coming out of night clubs as well. All seemed drugged by the powerful beat that they had been listening and dancing to for an entire night. People even walked in that rhythm — almost as if they were marching! “Mary, this scares me... people start and stop walking at beat intervals, and change direction and they do not act in between beats at all, except for keeping on moving!” “No problem, Jay, they are used to it. Besides, I’ve heard that everybody walks like that, everywhere in town. I heard it from a neighbour that went to Dallas a few days ago.”

Jay had a bad feeling. When they came to a non-functioning traffic light he could see why: The lights, that did not signal at the trusted exact intervals anymore, confused the people. Some refused to cross, while others took no notice of the wait signal anymore and just crossed without looking. This was going to cause accidents! Cars started to pull up from the parking lot, and approached the light. None of the street crossing people took notice of the cars! Jay jumped in, and started to guide the people over the street, as if he were a police man.

It took some while before Jay and Mary thought traffic had slowed down enough to make it safe to return to their car. While driving home, Mary openly showed her pride of Jay. “Jay, you did so well... You avoided some serious accidents out there!” Mary complimented him. “Maybe, but what had gotten into those people? Is that what the rhythm does to people? Does it stop them from thinking? In that case, I can understand why it is so popular! People are drugged, they get an easy way out of taking their responsibilities. No wonder they call us Jazzers; to them we must seem like utterly creative people!” Mary took no notice of Jay’s complaints, and turned on a radio station with Woop-Aloop.

* * * * *

T heir lives changed with the new synchronization at home. Mary took over the standard beat in her life, while Jay was a little more reluctant to do so. “You’re just like Simon. You guys are just too old to accept new ideas.” Then, continuing: “Talking about the devil... there he comes.” Simon indeed entered, with a shocked look on his face. “Simon, what is wrong with you? Are you al-right?” Mary immediately demanded. “I am fine, thanks, but have you heard the news? Austin is also raising the power of their synchronizing beacons! Same reason: to get more money. This may go on for a while. It means war between Austin and Dallas — and we are the catch they’re fishing at!”

“Maybe you need a drink, Simon? You seem pretty excited.” “No thanks, I want to think clearly — maybe we, being Jazzers and all, are the ones to prevent this war from happening.” Jay agreed. “Yes, let’s try to figure out a solution. Having seen the people in Dallas, I fear we are the only ones that can still think straight and prevent worse situations. We’ve been to Dallas yesterday. The situation is horrible, nobody even realizes that the rhythm has taken control of their entire lives. They’re sinking deeper and deeper, and are not capable of doing anything out of sync. We should make end this entire rhythm story!” Mary touched his hand softly, to show she felt the same way. She was now getting convinced that the synchronization was going to bring more pain than pleasure. The problem now was to come up with some tactics.

Mary sighed “If only we could get into these broadcasting stations we could simply turn them off. That would solve the problem, right?” “Yes, of course,” Jay responded, “but how to avoid that people keep on running around and synchronize with each other? Their rhythmical sense is strongly developed by now, and they should be capable of maintaining their rhythm together. If only we could get them away from their rhythm.” For a moment, they were silent and full of sorrow, but then Mary started laughing out loudly. She even got hick-ups when she explained that she was thinking of those Russian dances, where the rhythm would go faster and faster, and the dancers would keep kicking their legs. She had once tried to do this as a little kid, and already halfway the speedup she had simply fallen on the ground, incapable of moving to such fast rhythms.

Simon was in no mood for such funny anecdotes, but Jay compensated his lack of enthousiasm with a wild idea: “Just imagine that we can increase the speed of the rhythm... then all devices would blink faster and faster, and people would take over these faster rhythms... if we do it slowly enough at least... and at a certain point they simply cannot follow it anymore. Then they’ll be confused, but also they have been freed from the rhythms without a chance to maintain the beats! But how can we get access to the broadcast stations? Colonel?”

Simon got enthousiastic too. “Leave that to me. We’re Jazzers, remember? We are capable of improvising. We’ll use our rhythmical improvisation as our main strength and gain access to the broadcasting stations. Since they expect no attack, we should have no problems.” “But,” Mary added, “there will still be signals from other cities. If Dallas has no beacons left, maybe the devices will pick up the Austin signals, when these are the only ones left?” The colonel nodded. “You’re right, Mary, we need to make sure it happens on a large scale... I’ll have to get in touch with some old buddies from the army. They’re living all over the states, and several owe me one. I should be able to persuade them to join me in this battle against the worst enemy one can ever imagine — ourselves!”

Despite of his age, Simon sure was fit and ready to go. It had taken him only a week to organize the take-over of the beacons throughout the country. The time of action had been set to 4 AM, so Jay was still yawning. Mary on the other hand, was as awake as the colonel, simply because she was too tense. She felt like a burgler, standing in the bushes until it was time to take action.

The colonel signalled that they should start their carefully planned operation. There was a time scheduled for everything, so that the whole country would act the same at the same time. The job itself was not too hard, since all that had to be done was overrule some white collar workers inside the beacon.

When they approached the building, Mary staggered. There was a porter standing by the door! “How are we going to get past him?” she asked the colonel. “That’s simple, Mary,” was his well-prepared answer, “we will approach him in the rhythm of the city people. Only when we are close enough we will abuse the fact that we are Jazzers: We will make a sudden move halfway a beat and slip by him.” And indeed, this worked. Before the porter knew what had happened he was locked out of his own building!

They took the elevator up to the top floor of the building, where the beacon’s signals should be generated according to the information that Simon had somehow obtained — the old devil refused to explain where the information came from! Most important was, the information turned out accurate. There was nobody present in the beacon’s rhythm-generator room. That was easy.

“Well done, folks. Now, all that we need to do is wait until the other beacons have been taken over — that should be in nine minutes from now.” They waited during these minutes, and started increasing the beat rhythms in the speed that would be the same for the whole country. Most important was to make this change very, very slowly, to avoid shock reaction by the people in the streets. Mary looked out a small window at one side of the beacon room and started to laugh. “Look outside, it’s crazy!” Jay joined her, grinning “Laurel and Hardy couldn’t have done it any better.” It was getting funnier and funnier while the speed kept increasing. People just fell on the ground and sat there, looking totally helpless now their drug had gone, now they had to think for themselves a little more again. By noon, the last people started to loose the rhythm, simply because it was going too fast.

“Well, Simon,” Mary started, “you did a splendid job, those people don’t know what...” and then, surprised, “What?” Some of the people had climbed up again, and had simply started moving on half the broadcasted rhythm. That was a surprise, even to the colonel. Musicality was the last thing they had expected to be in their way!

The colonel started making nervous phone calls, to find out the thing he already feared: the same effect was taking place in all other cities as well. This was a disaster. But Jay’s technical instinct helped them out: “Why don’t we stay here and keep on increasing the rhythm?” “Because then they’ll respond to a quart of the rhythm, and then to an eighth, and so forth. They would always get around it!” “No they won’t,” Jay brought in, “at a certain point the result is not a rhythm anymore, but it goes so fast that it is simply a tone, and nobody can separate that out anymore.” “Splendid idea, Jay!” the colonel shouted. He started phoning it around, and the action was taken.

It took them three more days to make sure that the beacon’s frequency was so high up that nobody could separate out the beats anymore. As a matter of fact, a tone was hanging all over the city, and this was no problem to anyone. People liked the low tone, and started humming it to themselves. Then, more and more people joined, and what used to be a staccato beat that hung all over the city, had now turned into a low-pitched hum, that gave a pleasant, welcoming feeling and, most importantly, allowed pedestrians and traffic to take a normal flow. The accidents due to the synchronizing mania had come to an end.

“Can we just leave this room unguarded, Simon?” Jay wanted to know, “They might as well turn the old rhythm back on again.” “Don’t worry, Jay. They seem to be so much happier now, the tone is more pleasant to them as a reference than any beat can ever be. To be honest, I even feel the urge to hum along with it.” And indeed, Mary and even Jay caught themselves trying to pick up the tone too.

On their way home they saw happy people, singing and sometimes even dancing. They had done a good thing and were happy. Now that the weight of their immense responsibility fell off of their shoulders, they started feeling that they hadn’t seen a proper bed for several days, and they went home to compensate for that.

Mary was the first one to wake up. “Jay, are you already awake?” she inquired. He muttered a bit, but came to his senses when she kissed him sensually on the mouth. For once in his life, Jay woke up in a happy mood. And it only got better during the minutes that followed.

An hour later they had taken a shower, and were eating their breakfast. “Funny, isn’t it? We took the rhythm from other people, but now our rhythm of life is also messed up — eating our breakfasts at 3 PM.” Jay grinned. “Maybe we should go outside and see if the Jazzers in our neighbourhood are feeling better now. After all, they do now receive the broadcasts of tones, and that means they can live in the same, happy way the people in the cities can.”

Their happy mood vanished when they came outside. The neighbours, who had heard from Simon’s wife what they had done, did not give them the hero treat that they had deserved. On the contrary: “Ah, are you done with your little game? Maybe you should have left things the way they were. Then at least we could live our lives happily!” Mary looked at Jay, but he was as puzzled as her when he asked “What is wrong now, then? Don’t you hear soft tones that make you want to hum?” “Sure we do. And so do others. The only problem is that not everyone hums to the same beacon, and the tones they produce are not quite the same. Oh, sure they’re almost the same, but that’s what makes this humming of people so irritating: they hum out of tune!” Instead of getting upset with those ungrateful people, Mary took a deep breath and asked them for any ideas to solve this problem. This positive approach cooled down the neighbours, but they did not produce any useful solution.

W hen Simon came over to Jay and Mary’s, he was as surprised as they had been to hear about the reactions of their neighbours. “What did we do wrong? Should we have given all beacons the exact same frequency? And is it worthwhile to make that happen with another action? And...” Jay interrupted Simon’s train of thoughts. “It would certainly be worthwhile, having seen the agony, but it would not work. Even if you manage to get all beacons to send out the exact same frequency, then still there would be places where there is a dip in the hum sent out. A bit like anti-noise, actually, the beacons will reverse each others effects locally and leave no resulting hum.” Mary got the message. “So, when my washer/dryer would not work synchronously in the left side of the bath room, I would be forced to move it to the right side? That would be ridiculous!”

The mood was getting a little distressed. “If this is what you get for sticking out your neck, I don’t think I can find the energy to solve this problem. Gratitude would have been the least.” That was Simon, who had spent a great deal of time to the planning of the operation. “Oh, Simon, you did so well, it really hasn’t been your fault.” Jay sighed. He knew there was no chance that they would ever solve the new problem tonight. Not in this mood. Simon showed his dismay rather openly, but Jay felt the same way, and Mary probably tried to seem happier than she truly felt. After an hour of producing little more than an occasional deep sigh, they wished each other good-night and Simon went to his wife to find some comfort.

Jay was the easy-sleeping kind. He managed to fall asleep the instant he hit his pillow. Mary sighed too, now. She knew it would take her forever to fall asleep, given the problems they were facing. She knew she could find a solution, if only... well, probably if she spent all night worrying about it. And so she did, although not voluntarily. Around 2 AM she suddenly sat straight up in bed. “Jay, I’ve got the solution!” “Hmmm??? Yes, sure... go back to sleep, dear.” But Mary would not listen. She pushed Jay around until he had to wake up. “Honey, what is going on? What time... it’s not even close to seven o’clock! Why do you wake me up?” And when he saw the enthusiasm in her eyes: “Oh no, not now, please. I’ve got a headache...” “No, stupid! That’s not it! I solved the problem!” “Problem? Oh, with the tones you mean?” Jay got a little more awake now that his technical interest was raised. “Sure, all we have to do is make the tones in every overlapping region a harmonious chord! It will take some puzzling, but we can do it! Come on, out of bed!” The blankets were lifted, and the nights were cold enough to freshen Jay into an awaken state. They went down.

“Jay, get your map of North America, and a glass, will you?” “With milk?”

“Nothing in it, I need it for drawing circles!” Mary took the glass and a pencil, and started drawing circles on the map to represent the beacons, which gave a good idea of their overlapping areas. She started writing notes on another piece of paper franticly. “Can’t you see what we need to do, Jay? We need to compose music! Not just a normal tune, but a very special one... It should be right whatever your traveling line through the US is... A two-dimensional composition! Now, what music style and chord scheme shall we use?” Jay started having fun now that he understood what Mary was doing. If ever Einstein and Mozart had known each other... no, even they probably wouldn’t have thought of a multi-dimensional piece of music!

The rest of the plan was straightforward. The actions of changing the beacon’s settings were executed a week later, like they had done it before. Simon once again turned out to be an excellent organizer: He arranged a soldier to enter each of the beacons, and a musician on every area where the broadcasts from beacons overlapped, walky-talkies, and so on. The final step was made in making America happy.

This time, when they came home, their neighbourhood was all cheerful. No wonder, Mary had insisted upon a major tone scale. The television broadcasted live interviews with important businessmen who had come to enjoy traveling so much more now they swiftly were subjected to sequences of tones that formed a pleasant melody - different for all the different paths they traveled, but always nice, harmonious music.

The neighbourhood in which Jay and Mary lived was at least as cheerful as the people in the cities; they had combinations of tones to listen to. And now that all was in sync, people would hum along and feel encouraged to mix in with their own tone, resulting in harmony.

F inally, after all the American people had gone through, they had created a life in which they felt pleasant. As long as there were the hums, there were smiling people, and an overall attitude of friendliness spread over the country. Crime rates dropped suddenly, and people became more willing towards others, regardless of who or what they were.

People lived in harmony. The more musical ones, soon followed by less daring people, started to improvise on the scale, singing their songs, and continuing it when they were on the road, traveling from beacon to beacon. The hum got into a hymn, getting people higher than high of pleasure.

Then, on a thursday afternoon, a white American entered China Town, San Francisco... What were those noises? Autsch, another tone scale!